OpenStudy (konradzuse):

find the integral... 17arctan(x) dx

OpenStudy (anonymous):

pull out the 17 and then use your integration table

OpenStudy (konradzuse):

\[17[\frac{1}{1 + x^2} ]+ c\]

OpenStudy (konradzuse):

\[\frac{17}{1 + x^2} + c\]

OpenStudy (konradzuse):

I think this is correct? :P

OpenStudy (konradzuse):

hmm guess not... Then would it be 17 on the bottom also?

OpenStudy (anonymous):

all you have to do is prove that arctan is the derivative of what you have

OpenStudy (anonymous):

or you can do byparts if you want since that is how it's derived i believe

OpenStudy (konradzuse):

I didn't think we needed by parts for this....

OpenStudy (anonymous):

it's how it's derived i believe let me see

OpenStudy (konradzuse):

It seems like it though, wolfram has a weird answer like that so I guess so...

OpenStudy (konradzuse):

so u = arctan(x) \[du = \frac{1}{1 + x^2}\] dv = dx v = x

OpenStudy (konradzuse):

I can pul the 17 out right? or would it be v = 17x?

OpenStudy (anonymous):

dv=dx v=x u=arctan(x) \[du=\frac{1}{1+x^2}\] \[\int udv=uv-\int vdu\]

OpenStudy (anonymous):

you can pul it out it's a constant

OpenStudy (anonymous):

just make sure it multiplies everything

OpenStudy (konradzuse):

\[xarctan(x) - \int\limits \frac{x}{1+x^2} dx\]

OpenStudy (anonymous):

correct now can yu do the second intergral

OpenStudy (konradzuse):

\[17[xarctan(x) - \int\limits \frac{x}{1+x^2} dx]\]

OpenStudy (konradzuse):

so in this case u = 1+x^2 du = x \[\int\limits \frac{du}{u} \]

OpenStudy (konradzuse):

\[17xarctan(x) - 17\log(u) + c\]

OpenStudy (turingtest):

du=2xdx

OpenStudy (konradzuse):

\[17xarctan(x) - 17\log(1+x^2) + c\]

OpenStudy (konradzuse):

woopsies :P

OpenStudy (anonymous):

yeah

OpenStudy (konradzuse):

\[17xarctan(x) - \frac{17}{2}\log(1+x^2) + c\]

OpenStudy (konradzuse):

sorry huge headache and doing this too fast :P.

OpenStudy (konradzuse):

so apparently that's not right hmmph.

OpenStudy (anonymous):

that's not the right answer?

OpenStudy (konradzuse):

grr it wanted LN WTF!@@!~@~

OpenStudy (anonymous):

that's a bummer ask your teacher about it

OpenStudy (konradzuse):

It's all good, still got it right, but still booo!

OpenStudy (konradzuse):

I keep forgetting about LN...

OpenStudy (diyadiya):

Can i ask a question why isn't 17/(1+x^2)+C the answer ? ?

OpenStudy (konradzuse):

That's what I wanted to know lol :P.

OpenStudy (konradzuse):

I guess because that's only the derivative...

OpenStudy (anonymous):

beause the derivative of arctan(x) is u'/1+u^2

OpenStudy (anonymous):

you'd have to have something such as du/a^2+u^2 to do this

OpenStudy (konradzuse):

ugly...

OpenStudy (diyadiya):

@Outkast3r09 OMG!! Yes Thanksss :D

OpenStudy (konradzuse):

yupp...

OpenStudy (konradzuse):

the answer would be an arctan...

OpenStudy (diyadiya):

Uh-oh!!! my brain stopped working lol

OpenStudy (konradzuse):

Same!

OpenStudy (anonymous):

if it was like that... the parts method would have brought us back to it eventually depending on whether we needed to simplify or not